OpenStudy (anonymous):

Find dw/dt by using the chain rule approximation and convert w to a function of t before differentiating. w=cos(x-y),x=t^2,y=1

OpenStudy (anonymous):

I just want to know if I'm doing this right: \[dw/dt(dx.dt)+dw/dy(dy/dt)=-\sin(x-y)(2t)-0\] Plugging in both x and y: \[-\sin(t ^{2}-1)(2t)\]

OpenStudy (anonymous):

oh and from there my answer for dw/dt is -2tsin(t^2-1)

OpenStudy (anonymous):

Now doing the second part I get: \[w=\cos(t^2-1)=-2tsin(t^2-1)\] is this right?

OpenStudy (anonymous):

\[dw = (dw/dx)dx + (dw/dy)dy = -sen(x-y)dx +sen((x-y)dy\] that would be de dw now: dx = 2t, dy = 0 pluging in: \[-sen(t ^{2}-1)2t\] this would be dw/dt

OpenStudy (anonymous):

ok thanks.